OpenStudy (arihangdu):

In an experiment, a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. Find a quadratic model for the data and use it to estimate the population of bateria at 9 hours.

OpenStudy (arihangdu):

Please help me to solve the problems.

OpenStudy (arihangdu):

With step by step instructions.

OpenStudy (anonymous):

Let the equation be : at^2 + bt + c = P, where t = time (hrs), P = population (1000's). When t = 1, P = 3.03. When t = 2, P = 1.72. When t = 3, P = 1.17. Substitute these into the equation to obtain these 3 simultaneous equations : a + b + c = 3.03 4a + 2b + c = 1.72 9a + 3b = c = 1.17 Solving gives : a = 0.38, b = -2.45, c = 5.1. The equation is therefore, P = 0.38t^2 - 2.45t + 5.1 Testing with t = 0 to 6 gives the population values as provided, so it seems to be a valid model. At t = 9 hrs, P = 0.38*9^2 - 2.45*9 + 5.1 = 13.83.

OpenStudy (arihangdu):

How you get the solving directly ?

OpenStudy (anonymous):

#ilovegoogle

OpenStudy (arihangdu):

I know that much from google but I need steps as well

OpenStudy (anonymous):

Okay give me a second

OpenStudy (arihangdu):

Alright go head

OpenStudy (arihangdu):

How to solve this with steps by steps please help me will get madels and fan